I still don't get this!
Find the measures of the sides of each triangle:
The ratio of the measures of 3 sides of a triangle is 8:7:5. Its perimeter is 240 ft.
Can someone help me??
Thanks

Answer by Alan3354(69443)

8 + 7 + 5 = 20
240/20 = 12
8*12 = 96
7*12 = 84
5*12 = 60